OpenStudy (brittany16):

Solve for the variable. p - 1 = - 54

OpenStudy (anonymous):

add 1

OpenStudy (anonymous):

p=-52+1 p=-51

jimthompson5910 (jim_thompson5910):

p - 1 = - 54 p - 1+1 = - 54+1 p = - 53

OpenStudy (anonymous):

p=-53